OpenStudy (anonymous):

PLEASE HELP PLEASEEEEE!!!!! =*( sqrt 756 How do i do big numbers? Is there a way i can break this down easily?

OpenStudy (anonymous):

\[\sqrt{756}\] right?

OpenStudy (mokeira):

\[\sqrt{756}=\sqrt{36\times 21}\] basically, find two numbers that when you multiply give you 756, but ensure one number has a root root of 36 is 6 so \[6\sqrt{21}\]

OpenStudy (anonymous):

wait what?

OpenStudy (anonymous):

i need a fast method i cant just sit there at a test workin on this

OpenStudy (mokeira):

i have no other method

OpenStudy (anonymous):

oh use the squares hmmm

OpenStudy (anonymous):

Do you have a calculator with you?

OpenStudy (anonymous):

It's about 27. 5

OpenStudy (anonymous):

27.4954541697 to be exact

OpenStudy (dumbcow):

fastest way is keep dividing by 4 or 9 or whatever perfect square goes into it until you cant divide it any more 756/4 = 189 189/9 = 21 --> sqrt4 * sqrt9 * sqrt21 = 6sqrt21